OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. A. 18° B. 36° C. 54° D. 58°

OpenStudy (anonymous):

36

OpenStudy (anonymous):

x+y=x-y+72 2y=72 y=36